OpenStudy (anonymous):

Which of the following best represents the average rate at which a person can quickly walk? 3 steps per hour 3 steps per second 30 steps per minute 30 steps per second

OpenStudy (anonymous):

First you need to add all of so 3+3=6 30+30=60 so now let's add 60+6= ? and after you have solved that you need to divide then multiply and you will have your answer !!!!!!!!!!!! :)

OpenStudy (anonymous):

so, 3 steps per second, correct? @badgirlz11